OpenStudy (anonymous):

Janine made a cylindrical vase in which the sum of the lateral area and area of one base was about 3000 square centimeters. The vase had a height of 50 centimeters. Find the radius of the vase. Explain the method you would use to find the radius.

OpenStudy (anonymous):

3000= (h* l) + pi * r^2 l=2*pi*r 3000= (h * 2*pi*r) + pi * r^2 3000 = (50 * 2*pi*r)+ pi * r^2 3000=50pir + pir^2 what do I do after this?

sam (.sam.):

\[\Large A_{lateral}+A_{bottom}=3000 \\ \\ \Large 2\pi rh+\pi r^2=3000 \\ \\ \text{When} ~~h=50, r=?\]

sam (.sam.):

Yep you're in the right track, you need to factor that, use quadratic equation

sam (.sam.):

\[\pi r^2+50\pi r-3000=0 \\ \\ r=\frac{-50\pi \pm \sqrt{2500\pi ^2-4(\pi)(-3000)}}{2\pi}\]

OpenStudy (anonymous):

How do i simplify that?

sam (.sam.):

\[r=\frac{-50\pi \pm \sqrt{2500\pi^2+12000\pi}}{2\pi} \\ \\ r=-25 \pm \frac{\sqrt{2500\pi^2+12000\pi}}{2\pi} \\ \\ r=-25 \pm \frac{\sqrt{2500(3.142)^2+12000(3.142)}}{2(3.142)} \\ \\ r=-25 \pm 39.75 \\ \\ r=-25+39.75, ~~~r=-25-39.75 \\ \\ r=14.74, ~~~r=-64.75\]

sam (.sam.):

But since 'r' cannot be negative, then it is r=14.74cm

OpenStudy (anonymous):

Thank you !

sam (.sam.):

Wait hold on Your answer from 3000 = (50 * 2*pi*r)+ pi * r^2 Should get you 3000 = (100*pi*r)+ pi * r^2

sam (.sam.):

\[\pi r^2+100 \pi r=3000\]

sam (.sam.):

\[r=\frac{-100\pi \pm \sqrt{10000\pi ^2-4(\pi)(-3000)}}{2\pi} \\ \\ r=8.77cm\]

OpenStudy (anonymous):

How do you get so many different answers for "r"?

sam (.sam.):

The 14.74 is a mistake, Your answer from 3000 = (50 * 2*pi*r)+ pi * r^2 Should get you 3000 = (100*pi*r)+ pi * r^2 Then that should give you r=8.77cm

OpenStudy (anonymous):

Alright thank you!

OpenStudy (anonymous):

would the units be squared?

sam (.sam.):

no because it is length, not area or volume

sam (.sam.):

It's just cm

OpenStudy (anonymous):

thanks again.